Please do not delete this.
Suppose a baseball is shot up from the ground straight up with an initial velocity of 32 feet per second. A function can be created by expressing distance above the ground, s, as a function of time, t. This function is s = -16t2 + v0t + s0
• 16 represents ½g, the gravitational pull due to gravity (measured in feet per second 2).
• v0 is the initial velocity (how hard do you throw the object, measured in feet per second).
• s0 is the initial distance above ground (in feet). If you are standing on the ground, then s0 = 0.

How long will it take to hit the ground?

Answer by stanbon(75887)

Suppose a baseball is shot up from the ground straight up with an initial velocity of 32 feet per second. A function can be created by expressing distance above the ground, s, as a function of time, t. This function is s = -16t2 + v0t + s0
• 16 represents ½g, the gravitational pull due to gravity (measured in feet per second 2).
• v0 is the initial velocity (how hard do you throw the object, measured in feet per second).
• s0 is the initial distance above ground (in feet). If you are standing on the ground, then s0 = 0.
---------------------
s(t) = -16t^2+32t+0
------------------------
How long will it take to hit the ground?
When it hits the ground s(t)=0
solve for "t":
-16t^2+32t=0
-16t(t-2) = 0
t = 0 or t=2
It will start at the ground and be back on the ground after 2 seconds.
====================
Cheers
Stan H.